OpenStudy (anonymous):

Solve the equation 2 over 3 k + 6 = -12 for k. Please show your work and solution in the space provided.

OpenStudy (shamil98):

\[\frac{ 2 }{ 3k+6 } = -12\]

OpenStudy (anonymous):

\[\frac{ 2 }{3k+6}=-12\] Did I rewrite that correctly?

OpenStudy (shamil98):

Multiply 3k + 6 to both sides \[2 = -12(3k+6)\]

OpenStudy (shamil98):

Distribute the -12 \[2 = -36k-72\] Then add 72 to both sides \[74 =-36k\]

OpenStudy (shamil98):

\[\frac{ -74 }{ 36 } = k\]

OpenStudy (shamil98):

Simplify \[\frac{ -37 }{ 18 }\]